Please help me solve this word problem. A total of $39,000 is invested in two bonds that pay 5.75% and 6.25% interest. The investor wants to receive $2400 interest each year. How much should be invested in each account?
Answer by stanbon(75887)

Please help me solve this word problem. A total of $39,000 is invested in two bonds that pay 5.75% and 6.25% interest. The investor wants to receive $2400 interest each year. How much should be invested in each account?
----------------
Equation::
interest + interest = interest
0.0575x + (0.0625)(39,000-x) = 2400
----------------------
Multiply thru by 10,000 to get:
575x + 625*39,000 - 625x = 24000000
-----------------------
-50x = -375000
x = $7500 (amt. invested at 5.75%
39,000-7500 = $31.500 (amt. invested at 6.25%
---------------------------
Cheers,
Stan H.
==================